For a group of 100 men subjected to a stress situation, the mean number of heart beats per minute was 130, and the standard deviation was 5. Find the 90% confidence interval of the true mean.
1
Expert's answer
2020-10-26T19:49:02-0400

Given that,


population size n = 100

population mean "\bar x" = 130

Standard deviation "\sigma" = 5


Since we have to calculate the Confidence Interval C.I. at 90% the value of Z"\alpha/2" = Z0.05 since its a 2 tailed test.


From the standard normal distribution table we get the value of Z"\alpha/2" = Z0.05 = 1.645


Formula for C.I = "\bar x\pm\ Z_{\alpha/2} *\frac {\sigma}{\sqrt n}\ =\ 130\ \pm\ 1.645*\frac {5}{\sqrt{100}}"


Hence we get,


90% Confidence Interval = (129.1775, 130.8225)